# Beyond Subspace Isolation: Many-to-Many Transformer for Light Field Image Super-resolution

### v1

@@ Abstract
The effective extraction of spatial-angular features plays a crucial role in light field image super-resolution (LFSR) tasks, and the introduction of convolution and Transformers leads to significant improvement in this area. Nevertheless, due to the large 4D data volume of light field images, many existing methods opted to decompose the data into a number of lower-dimensional subspaces and perform Transformers in each sub-space individually. As a side effect, these methods inadvertently restrict the self-attention mechanisms to a One-to-One scheme accessing only a limited subset of LF data, explicitly preventing comprehensive optimization on all spatial and angular cues. In this paper, we identify this limitation as subspace isolation and introduce a novel Many-to-Many Transformer (M2MT) to address it. M2MT aggregates angular information in the spatial subspace before performing the self-attention mechanism. It enables complete access to all information across all sub-aperture images (SAIs) in a light field image. Consequently, M2MT is enabled to comprehensively capture long-range correlation dependencies. With M2MT as the pivotal component, we develop a simple yet effective M2MT network for LFSR. Our experimental results demonstrate that M2MT achieves state-of-the-art performance across various public datasets. We further conduct in-depth analysis using local attribution maps (LAM) to obtain visual interpretability, and the results validate that M2MT is empowered with a truly non-local context in both spatial and angular subspaces to mitigate subspace isolation and acquire effective spatial-angular representation.
@@ I Introduction
Light field (LF) images, unlike regular images captured by monocular cameras, provide richer information by capturing light rays from multiple angular directions in a single shot. This unique characteristic has paved the way for substantial progress in various computer vision applications where conventional cameras have shown limited efficacy, e.g., material recognition [1, 2], depth estimation [3, 4, 5, 6, 7], salient object detection under complex scenarios [8, 9, 10], microsopy [11, 12, 13], and anti-spoof face recognition [14, 15]. By simultaneously capturing multiple sub-aperture images (SAIs, or views), LF technology enables a rich and interactive viewing experience. Users can freely explore and interact with the virtual environments, changing perspectives and moving within them. Therefore, LF technology has become a cornerstone of VR applications, enhancing user engagement and immersion.
@@
EPIT (DI: 5.9518) M2MT (DI: 27.2578)
(b) Local attribution maps of SAIs: Diffusion Index (DI) quantifies the extent of influential pixels.
Figure 1: Super-resolved results and local attribute maps (LAM) of the proposed M2MT-Net against the current state-of-the-art methods on the Perforated_metal_3 sample.
Capturing LF images necessitated self-built dense camera arrays [16, 17], which were prohibitively expensive and not ready for mainstream use. However, advancements of sophisticated LF cameras like Raytrix [18], Lytro Illum [19], and Google’s Light Field VR Camera [20] have gradually democratized LF imaging, making it accessible for both commercial and industrial applications. Despite this progress, LF cameras have long faced the challenges in striking a balance angular and spatial resolutions due to inherent limitations in sensor capabilities, often leading to lower spatial resolutions compared to traditional cameras.
Researchers have developed a number of possible solutions and they generally fall into two categories: light field image super-resolution (LFSR) [21, 22] and light field angular super-resolution (also known as light field view synthesis) [23, 24]. LFSR aims to upsample the spatial resolution of each SAI, effectively reconstructing the details within the views, which is the focus of this paper. On the other hand, light field angular super-resolution focuses on synthesizing additional SAIs to enhance the angular resolution of the light field. This paper primarily concentrates on LFSR.
@@
Figure 2: Illustration of available data in LF tensors used by existing LF Transformers under the One-to-One scheme and our proposed Many-to-Many Transformer. For the tensors, each color represents a SAI.
@@
Motivations. The advances in deep learning, particularly convolutional neural networks (CNNs) [25, 21, 22] and Vision Transformers (ViT) [26, 27, 28], have led to significant improvement in LFSR than traditional methods [29]. The success has also extended to LFSR by processing 4D LF images in 2D subspaces, such as spatial, angular [30, 31], or Epipolar Image (EPI) [32] subspaces. However, these methods predominantly suffer from subspace isolation, a defect causing sub-optimal performance.
@@
Specifically, this issue arises when adapting 2D operations to the 4D LF data, existing approaches have to compromise its complete access to the LF information. This is primarily because training networks directly on voluminous 4D LF data, e.g. 4D convolutions [21], demands a relatively large number of weights, which is prone to optimization difficulties and heavy computation. As a workaround, most previous methods decompose the 4D data into 2D subspaces such as spatial and angular subspaces, or EPI subspaces. In implementation, one typical practice is to temporarily reshape a 4D tensor to expose two operative dimensions as tokens while merging the other two dimensions with the batch dimension. This decomposition enables 2D operations to perform on 4D LF tensors, and in training, the optimization is conducted on the whole tensor. However, a significant limitation arises during inference. When inferring the value at a specific location, access to the two merged dimensions is confined to only one location at a time, rather than spanning the entire dimensionality. As a result, even if with non-local Transformers, the effective receptive field is virtually restricted to a local context within the operative subspaces, leading to an One-to-One scheme.
@@
For instance, consider the scenario depicted in Figure 2(a), where a single 4D LF tensor $I(u,v,x,y)\in\mathbb{R}^{U\times V\times W\times H\times C}$ is part of a batch $B$. Here, $(u,v,x,y)$ denotes a pixel’s spatial location $(x,y)$ and angular location (or SAI) $(u,v)$. By merging the angular subspace $(U,V)$ into the batch dimension, a 2D spatial Transformer in Figure 2(b) is enabled to operate on the flatten spatial subspace $(W,H)$ as tokens across SAIs. However, this merging effectively isolates the network’s forward propagation in the merged angular subspace $BUV$, as depicted by transparent blocks in the batch dimension. Consequently, the network is restricted to accessing only one SAI at a time during processing. This procedural constraint can be formally expressed as
$\begin{split}I_{2}(u,v,x,y)=F\cdot\{I_{1}(\bar{u},\bar{v},\bar{x},\bar{y})\}_{(\bar{u},\bar{v})=(u,v),(\bar{x},\bar{y})\in\mathbb{R}^{W\times H}}\end{split}$ (1)
@@
where $F$ is a 2D operation, which can be either convolution or Transformers, and $I_{1}$ and $I_{2}$ are the input and output LF tensors of the operation. Under this scheme, to obtain a complete LF tensor, Equation 1 is repeated $U\times V$ times, each effectively being a One-to-One operation mapping from a SAI in $I_{1}$ at a single angular location $(u,v)$ to a SAI in $I_{2}$ at the same isolated angular location $(u,v)$ in the output. However, the ideal processing would instead use all SAIs to inform the computation loosening the constraint $(\bar{u},\bar{v})=(u,v)$, resulting in:
$\begin{split}I_{2}(u,v,x,y)=F\cdot\{I_{1}(\bar{u},\bar{v},\bar{x},\bar{y})\}_{(\bar{u},\bar{v},\bar{x},\bar{y})\in\mathbb{R}^{U\times V\times W\times H}}\end{split}$ (2)
@@
Subspace isolation is not unique to spatial Transformers and extends to other forms of data decomposition under the One-to-One scheme. For example, an angular Transformer is limited to accessing only one pixel across SAIs as depicted in Figure 2(c). Similarly, an EPI Transformer can only access a slice of two dimensions in the EPI subspace merged with the batch dimension, as shown in Figure 2(d). These constraints, inherent in the One-to-One operational scheme, significantly impede the ability of existing models to fully exploit the spatial and angular cues available in LF data, resulting in an incomplete spatial-angular representation.
@@
Contributions. To address this issue, in this paper, we propose the novel Many-to-Many Transformer (M2MT), a new scheme to achieve the goal of comprehensive data integration outlined in Equation 2 and alleviate the isolation. The M2MT method begins by constructing correlation embeddings in the angular subspace. It then applies a self-attention mechanism to model long-range dependencies within the spatial subspace. This innovative approach allows the M2MT to access all the spatial and angular cues present in an LF image during each step of data propagation, thereby facilitating the creation of a holistic spatial-angular representation with a truly non-local context.
@@
With M2MT as a foundational component, we present a simple yet effective network, M2MT-Net, incorporating M2MT in the spatial subspace and vanilla Transformers in the angular subspace. Through extensive experimental evaluation, we showcase M2MT-Net’s outstanding performance, establishing it as a new state-of-the-art for LFSR.
@@
Furthering the research, we delve into a series of analysis to discover the mechanisms behind its success. Particularly, by leveraging the technique of local attribution maps (LAM) [33], which visualize influential pixels, to gain interpretability of neural networks. Figure 1 reveals that M2MT utilizes more pixels across broader SAIs than the current state-of-the-art methods like EPIT [32]. This observation substantiates the efficacy of M2MT, which mitigates the limitation of subspace isolation, simultaneously preserving more high-frequency cues in the spatial subspace and establishing a richer and non-local interplay of SAI dependencies in the angular subspace.
@@
    The limitation of subspace isolation is identified, which predominantly impedes the performance of existing LF image processing methods by restricting data access when decomposing a 4D operation into a few One-to-One 2D operations.
@@
    To overcome the limitations imposed by the One-to-One scheme, we introduce the Many-to-Many Transformer (M2MT), which aggregates correlation information from the angular subspace and performs a self-attention mechanism in the spatial subspace. A simple-yet-effective network, M2MT-Net, is developed incorporating the proposed M2MT for LFSR.
@@
    Experimental results validate that M2MT-Net achieves state-of-the-art performance. In-depth analysis, especially the LAM analysis, is conducted to reveal its truly non-local context, utilizing all available information in a LF image for superior performance.
II Related Works
@@ II-B Light Field Image Super-resolution
Processing 4D LF data presents significant challenges in developing neural networks. The application of 4D convolutions is a straightforward solution but results in computationally heavy models, making both training and inference difficult. To alleviate this drawback, Wang et al. [1] introduced an interleaved filter as an approximation for light field material recognition. The filter decomposes a 4D decomposition into a spatial and angular convolution. They proved that comparable performance can be achieved by interleaving these two distinct convolutions. This decomposition scheme is adopted by Yoon et al. in LFCNN for LFSR [25]. LFCNN consists of a spatial sub-network for SAI processing and an angular sub-network composed of three branches to capture LF correlation in three different geometric directions. Yeung et al. [21] proposed a deep neural network consisting of a series of spatial-angular separable (SAS) convolution, akin to interleaved filters but trained in an end-to-end manner. Jin et al. [48] proposed an all-to-one framework, where each SAI is individually super-resolved using the other SAIs. A structure-aware loss is incorporated to preserve LF images’ inherent parallax structure. Wang et al. [49] introduced a network to extract spatial and angular features in separate branches and iteratively fuse them. Liu et al. [50] proposed a pyramid network with dilated convolutions to expand receptive fields in both spatial and angular subspaces. Chen et al. [51] incorporates the frequency domain and semantic prior and proposed a network to super-resolve both spatial and angular resolutions. Sun et al. [52] proposed an network with disparity-exploited and non-disparity branches to learn a compact spatial-angular representation. Hu et al.[53] proposed Decomposition Kernel Network (DKNet), which generalizes the decomposition scheme to comprehensively cover the spatial, angular and EPI subspaces. Further advancing the scheme, Wang et al. [22] proposed a disentangling mechanism to enhance the effectiveness. Different from the previous methods, some works resorts to non-deep-learning-based models [54, 55]. Some works explored plug-and-play strategies to boost the performance of existing methods, like the learning prior from single images [56] and the cut-and-blend data augmentation [57].
@@
In parallel to SISR, ViT has broadened the LFSR landscape. DPT [30] leveraged Transformers to learn image and gradient information among SAIs in horizontal and vertical sequences. LFT [31] drew parallels with the earlier decomposition scheme but employed Transformers in place of separable convolutions. To enable spatial Transformers to model both local and non-local dependencies, spatial features were locally unfolded into patches and subsequently processed through a linear layer for local embedding before self-attention. EPIT proposed to utilize Transformers in horizontal and vertical EPI subspaces. Liang et al. proposed EPIT [32] to further explore the use of Transformers in horizontal and vertical EPI subspaces.
@@
Despite these advancements, a common limitation predominantly persists across the most decomposition-based methods: subspace isolation, as elaborated in Section I. This limitation motivates the derivation of our work in this paper.
@@
Figure 3: Illustration of M2MT-Net and its components. (a) depicts the overview of M2MT. (b) and (c) illustrate the details of a M2MT Transformer and an angular Transformer. These two components constitute a Correlation Block in (d).

### v2

@@ Abstract
The effective extraction of spatial-angular features plays a crucial role in light field image super-resolution (LFSR) tasks, and the introduction of convolution and Transformers leads to significant improvement in this area. Nevertheless, due to the large 4D data volume of light field images, many existing methods opted to decompose the data into a number of lower-dimensional subspaces and perform Transformers in each sub-space individually. As a side effect, these methods inadvertently restrict the self-attention mechanisms to a One-to-One scheme accessing only a limited subset of LF data, explicitly preventing comprehensive optimization on all spatial and angular cues. In this paper, we identify this limitation as subspace isolation and introduce a novel Many-to-Many Transformer (M2MT) to address it. M2MT aggregates angular information in the spatial subspace before performing the self-attention mechanism. It enables complete access to all information across all sub-aperture images (SAIs) in a light field image. Consequently, M2MT is enabled to comprehensively capture long-range correlation dependencies. With M2MT as the foundational component, we develop a simple yet effective M2MT network for LFSR. Our experimental results demonstrate that M2MT achieves state-of-the-art performance across various public datasets, and it offers a favorable balance between model performance and efficiency, yielding higher-quality LFSR results with substantially lower demand for memory and computation. We further conduct in-depth analysis using local attribution maps (LAM) to obtain visual interpretability, and the results validate that M2MT is empowered with a truly non-local context in both spatial and angular subspaces to mitigate subspace isolation and acquire effective spatial-angular representation.
@@ I Introduction
Light field (LF) images, unlike regular images captured by monocular cameras, provide richer information by capturing light rays from multiple angular directions in a single shot. This unique characteristic has paved the way for substantial progress in various computer vision applications where conventional cameras have shown limited efficacy, e.g., material recognition [1, 2], depth estimation [3, 4, 5, 6, 7], salient object detection under complex scenarios [8, 9, 10], microscopy [11, 12, 13], and anti-spoof face recognition [14, 15]. By simultaneously capturing multiple sub-aperture images (SAIs, or views), LF technology enables a rich and interactive viewing experience. Users can freely explore and interact with the virtual environments, changing perspectives and moving within them. Therefore, LF technology has become a cornerstone of VR applications, enhancing user engagement and immersion.
@@
EPIT M2MT-Net
DI: 8.3895 DI: 25.1975
(b) Local attribution maps of SAIs. Diffusion Index (DI) quantifies the extent of influential pixels.
Figure 1: Super-resolution results and local attribute maps (LAM) of the proposed M2MT-Net against EPIT on the Perforated_metal_3 sample.
Capturing LF images necessitated self-built dense camera arrays [16, 17], which were prohibitively expensive and not ready for mainstream use. However, advancements in sophisticated LF cameras like Raytrix [18], Lytro Illum [19], and Google’s Light Field VR Camera [20] have gradually democratized LF imaging, making it accessible for both commercial and industrial applications. Despite this progress, LF cameras have long faced challenges in striking a balance of angular and spatial resolutions due to inherent limitations in sensor capabilities, often leading to lower spatial resolutions compared to traditional cameras.
Researchers have developed a number of possible solutions, and they generally fall into two categories: light field image super-resolution (LFSR) [21, 22, 23] and light field view synthesis (also known as light field reconstruction or light field angular super-resolution) [24, 25, 26, 27, 28, 29]. LFSR aims to enhance the spatial resolution of all SAIs, while light field view synthesis focuses on synthesizing additional SAIs to enhance the angular resolution of a light field image. Additionally, some works [30, 31, 32] have developed methods for joint super-resolution that enhance both spatial and angular resolutions simultaneously. In this paper, we primarily concentrate on LFSR.
@@
Figure 2: Illustration of accessible data in LF tensors used by existing LF Transformers under the One-to-One scheme and our proposed Many-to-Many Transformer. For the LF tensors, each color represents a SAI.
@@
Motivations. The advances in deep learning, particularly convolutional neural networks (CNNs) [33, 34] and Vision Transformers (ViT) [35, 36, 37], have led to significant improvement in LFSR than traditional methods [38]. Among them, most methods have opted to process 4D LF images in their 2D subspaces, such as spatial, angular [39, 40, 23, 22], or Epipolar Image (EPI) [41, 22] subspaces. However, these methods predominantly suffer from subspace isolation, a defect causing sub-optimal performance.
@@
Specifically, when adapting 2D operations to the 4D LF data, existing methods have to compromise their complete access to the LF information. This is primarily due to training networks directly on voluminous 4D LF data, e.g., 4D convolutions [21], demands a relatively large number of weights, which is prone to optimization difficulties and heavy computation. As a workaround, most previous methods decompose the 4D data into 2D subspaces such as spatial and angular subspaces, or EPI subspaces. In implementation, one typical practice is to temporarily reshape a 4D tensor to expose two operative dimensions while merging the other two dimensions with the batch dimension. This decomposition enables 2D operations to perform on 4D LF tensors, and in network training, the optimization is conducted on the whole tensor. However, a significant limitation arises during inference. When inferring the value at a specific location, access to the two merged dimensions is confined to only one location at a time rather than spanning the entire dimensionality. As a result, even with non-local Transformers, the effective receptive field is virtually restricted to a local context within the operative subspaces, leading to a local One-to-One scheme.
@@
For instance, considering the scenario depicted in Fig. 2(a), a single 4D LF tensor within a batch $B$ is defined as $I(u,v,x,y)\in\mathbb{R}^{U\times V\times W\times H\times C}$, where $U$ and $V$ denote the two angular dimensions, $W$ and $H$ denote the two spatial dimensions, and $C$ denotes the channel dimension. Here, $(u,v,x,y)$ denotes a pixel’s spatial location $(x,y)$ and angular location (or SAI) $(u,v)$. By merging the angular subspace $(U\times V)$ into the batch dimension $B$, a 2D spatial Transformer in Fig. 2(b) is enabled to operate on the flattened spatial subspace $(W\times H)$ as tokens across SAIs. However, this merging operation virtually isolates the network’s forward propagation within only a location in the merged angular subspace $BUV$. The accessible data in the batch dimension is depicted by the opaque block, while the isolated data is transparent. Consequently, the network’s receptive field is restricted to only one SAI at a time during inference. This procedural constraint can be formally expressed as
$\begin{split}I_{2}(u,v,x,y)=F_{O2O}\cdot\{I_{1}(\bar{u},\bar{v},\bar{x},\bar{y})\}_{(\bar{u},\bar{v})=(u,v),(\bar{x},\bar{y})\in\mathbb{R}^{W\times H}}\end{split}$ (1)
@@
where $F_{O2O}$ represents a One-to-One operation, which can be either convolution or Transformers, and $I_{1}$ and $I_{2}$ are the input and output LF tensors of the operation. Under this scheme, to obtain a complete LF tensor, Equation 1 must be repeated $U\times V$ times mapping from a SAI in $I_{1}$ at a single angular location $(\bar{u},\bar{v})$ to a SAI in $I_{2}$ at the same isolated angular location $(u,v)$ in the output. However, the ideal processing would instead use all SAIs to inform the calculation loosening the constraint $(\bar{u},\bar{v})=(u,v)$, resulting in a Many-to-Many operation $F_{M2M}$:
$\begin{split}I_{2}(u,v,x,y)=F_{M2M}\cdot\{I_{1}(\bar{u},\bar{v},\bar{x},\bar{y})\}_{(\bar{u},\bar{v},\bar{x},\bar{y})\in\mathbb{R}^{U\times V\times W\times H}}.\end{split}$ (2)
@@
Subspace isolation is not unique to spatial Transformers and extends to other forms of data decomposition under the One-to-One scheme. For example, an angular Transformer is limited to accessing only one pixel across SAIs, as depicted in Fig. 2(c). Similarly, an EPI Transformer can only process a two-dimensional slice within the EPI subspace at one step. Specifically, it can handle a horizontal EPI $\mathbb{R}\in R^{U\times W}$ or a vertical EPI $\mathbb{R}\in R^{V\times H}$ as illustrated in Fig. 2(d), while the remaining slices stay isolated in the batch dimension. These constraints, inherent in the local One-to-One operational scheme, significantly impede the ability of existing models to fully exploit the spatial and angular cues available in LF data, resulting in an incomplete spatial-angular representation.
@@
Contributions. To address this issue, in this paper, we propose the novel Many-to-Many Transformer (M2MT), a new scheme to achieve the goal of comprehensive data integration outlined in Equation 2 and alleviate the isolation. The M2MT method begins by constructing a correlation tensor in the angular subspace. It then applies a self-attention mechanism to model long-range dependencies within the spatial subspace. This innovative approach allows the M2MT to access all the spatial and angular cues present in an LF image during each step of data propagation, thereby facilitating the creation of a comprehensive spatial-angular representation with a truly non-local context.
@@
With M2MT as a foundational component, we present a simple yet effective network, M2MT-Net, incorporating M2MT in the spatial subspace and vanilla Transformers in the angular subspace. Through extensive experimental evaluation, we showcase M2MT-Net’s outstanding performance and an excellent performance-efficency balance, establishing it as a new state-of-the-art for LFSR.
@@
Furthering the research, we delve into a series of studies to discover the mechanisms behind its success. Particularly, by leveraging the technique of local attribution maps (LAM) [42], which visualize influential pixels, to gain interpretability of neural networks. Fig. 1 reveals that M2MT-Net utilizes more pixels across broader SAIs than the current state-of-the-art methods like EPIT [41]. This observation substantiates the efficacy of M2MT-Net, which mitigates the limitation of subspace isolation, simultaneously preserving more high-frequency cues in the spatial subspace and establishing a richer and non-local interplay of SAI dependencies in the angular subspace.
We also conducted an analysis using light field depth estimation to validate the angular consistency in the reconstructed results. The results demonstrate that M2MT-Net’s depth maps are sharper and more integrated, suggesting that, besides reconstructing more visual details, M2MT-Net effectively preserves the parallax structure across SAIs, enriching the realism of the resulting LF images.
@@
    We propose the Many-to-Many Transformer (M2MT), a novel approach integrating spatial and angular information in light field images. By constructing a correlation tensor in the angular subspace and applying a self-attention mechanism in the spatial subspace, M2MT addresses the subspace isolation prevalent in the previous methods by its truly non-local context.
@@
    We introduce M2MT-Net, which incorporates M2MT in the spatial subspace and vanilla Transformers in the angular subspace. Extensive experiments show that M2MT-Net sets a new state-of-the-art for LFSR in terms of performance. Furthermore, M2MT-Net strikes a compelling balance between model performance and efficiency, producing higher-quality LFSR results with substantially lower memory and computation requirements.
@@
    We provide insights into M2MT-Net’s effectiveness. The analysis of local attribution maps (LAM) is conducted to visualize influential pixels, showing that M2MT-Net utilizes more pixels across a broader range of sub-aperture images (SAIs) compared to existing methods. Additionally, our analysis of light field depth estimation reveals that M2MT-Net produces sharper and more integrated depth maps, which suggests that it preserves the parallax structure of LF images, enhancing its realism with better angular consistency.
II Related Work
@@ II-B Light Field Image Super-resolution
Processing 4D LF data presents significant challenges in developing neural networks. The application of 4D convolutions is a straightforward solution but results in computationally heavy models, making both training and inference difficult.
To alleviate this drawback, Farrugia et al. [54] proposed a framework incorporating optical flow and a deep CNN to reduce the angular subspace to construct a compact representation preserving angular consistency and subsequently restore the whole LF image. Wang et al. [1] introduced an interleaved filter as an approximation for light field material recognition. The filter decomposes a 4D convolution into a spatial convolution and an angular convolution. They proved that comparable performance can be achieved by interleaving these two distinct convolutions.
@@
This decomposition scheme was adopted by Yoon et al. in LFCNN for LFSR [55]. LFCNN consists of a spatial sub-network for SAI processing and an angular sub-network composed of three branches to capture LF correlation in three different geometric directions. Yeung et al. [21] proposed a deep neural network consisting of a series of spatial-angular separable (SAS) convolution, akin to interleaved filters but trained in an end-to-end manner. Jin et al. [56] proposed an all-to-one framework where each SAI is individually super-resolved using the other SAIs. A structure-aware loss is incorporated to preserve LF images’ inherent parallax structure. Wang et al. [57] introduced a network to extract spatial and angular features in separate branches and iteratively fuse them. Liu et al. [58] proposed a pyramid network with dilated convolutions to expand receptive fields in both spatial and angular subspaces. Chen et al. [59] incorporated the frequency domain and semantic prior and proposed a network to super-resolve both spatial and angular resolutions. Sun et al. [60] proposed a network with disparity-exploited and non-disparity branches to learn a compact spatial-angular representation.
Further advancing the scheme, Cheng et al. [27] proposed the concept of spatial-angular correlated convolution, extending the SAS scheme [21] to the EPI subspaces and forming spatial-angular versatile convolution (LFSSR-SAV). Hu et al.[61] proposed the Decomposition Kernel Network, which generalizes the decomposition scheme to comprehensively cover the spatial, angular and EPI subspaces. Wang et al. [22] proposed a disentangling mechanism to aggregate and enhance features from these subspaces. Duong et al. [62] combined the angular and spatial extractors with its proposed multi-orientation epipolar extractors to cover more aspects of LF images.
Different from the previous methods, some works resort to non-deep-learning-based models [63, 64]. Some works explored plug-and-play strategies to boost the performance of existing methods, like the learning prior from single images [65] and the cut-and-blend data augmentation [66].
@@
In parallel to SISR, ViT has broadened the LFSR landscape. DPT [39] leveraged Transformers to learn image and gradient information among SAIs in horizontal and vertical sequences. LFT [40] drew parallels with the earlier decomposition scheme but employed Transformers in place of separable convolutions. To enable spatial Transformers to model both local and non-local dependencies, the spatial features were locally unfolded into patches and subsequently processed through a linear layer into local embeddings before self-attention. Liang et al. proposed EPIT [41] to further explore the use of Transformers in horizontal and vertical EPI subspaces. To enhance the capability of spatial and angular Transformers, Cong et al. proposed a sub-sampling spatial modeling strategy and a multi-scale angular modeling strategy in their LF-DET [23].
@@
Despite these advancements, a common limitation predominantly persists across most decomposition-based methods: subspace isolation, as elaborated in Section I. This limitation motivates the derivation of our work in this paper.
@@
Figure 3: Illustration of M2MT-Net and its components. (a) depicts the overview of M2MT. (b) and (c) illustrate the details of a M2MT Transformer and an angular Transformer. These two components constitute a Correlation Block in (d). $\bigoplus$ represents the addition operation of a residual connection.